# From Specificity to Generality: Revisiting Generalizable Artifacts in Detecting Face Deepfakes

### v1

@@ 4.3 Ablation Study
The impact of method components.
To evaluate the effectiveness of each component in our method, we performed ablation studies on multiple datasets. In Table 6(a), we present the results of our ablation experiments on FIA-USA, Automatic Forgery-aware Feature Selection (AFFS), and Region-aware Contrastive Regularization (RCR), as well as the ablation of components in FIA-USA and RCR. The experimental results in the Table 6(a) underscore the efficacy of each component of our method. Notably, we report Video-Level AUC for traditional deepfake datasets and Frame-Level AUC for generative deepfake datasets.
The impact of model framework. Concurrently, we also examined the influence of various model architectures, as detailed in the Table 6(b). We assessed the performance implications of EfficientNet [71], Resnet [30], and network layers on model performance. To ensure a fair comparison with ResNet, we extended training by an additional 100 epochs to achieve a training loss comparable to that of EfficientNet.
Impact of hyperparameters. We also examined the effect of hyperparameters of the loss function on model performance, as detailed in the Table 6(c). Our analysis of the weights revealed that increasing the weight of $\lambda_{2}$ relative to the weights of cross entropy enhances model performance, whereas decreasing the weight of $\lambda_{3}$ relative to cross entropy also improves performance.
@@ 5 Visualizations
We employed GradCAM [95] to identify the regions activated for forgery detection. The visualization results depicted in Figure 4 demonstrate that our model precisely identifies the areas of deepfake forgery. Beyond visualizing tampered traditional deepfake images, we also conducted visualizations on images produced by state-of-the-art generation techniques.
@@ Appendix A Appendix
  • Sec. I: We provided more visual examples.
@@ Appendix D Details of FPN
As shown in the Figure 11, our FPN follows the framework proposed by Lin et al., using the output features of the 2nd, 3rd, 4th, and 5th layers of the backbone to construct an FPN, In our experimental setup, the number of feature map channels is 128 and 196, corresponding to EfficientNet [71], and Resnet [30], respectively.
@@ E.2 AFFS
Here, we will provide a more comprehensive explanation of the workflow of AFFS from a different perspective.
1) First, let’s consider multiple positive and negative sample pairs generated by FIA-USA, along with their corresponding masks $\{R_{n},F_{n},M_{n}\}_{n=1}^{N}$. The usual method is to directly construct an FPN using a set of features $f_{i}$ from a neural network $\phi$ as input, where $i$ represents the $i$-th layer of the $\phi$. We can see that this process is not related to FIA-USA, so it cannot fully unleash the potential of FIA-USA, and the constructed feature pyramid also intuitively lays the foundation for RCR.
@@
$\mathcal{F}$ represents normalization and scaling operations, $f_{i}^{k}\in\mathbb{R}^{W_{i}*H_{i}},k\in[0,C_{i}]$, represents the feature map of the $k-$th dimension of the $f_{i}$ During the training process, we only select the channels reserved for $f_{i}$ during the AFFS initialization process. After AFFS processing, the original features $f_{i}\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ will become $f_{i}^{\prime}\in\mathbb{R}^{W_{i}*H_{i}*m_{i}}$
@@ E.3 Visual Explanation of AFFS
Here we provide a set of visual examples of AFFS in Figure 13. The experiment was conducted on the 2-th layer of EfficientNetB4. Unlike the actual situation, in order to enhance readability, we only calculated $L_{AFFS}$ for a single epoch on a single pair of samples. And based on this, sort the 32 dimensions of the features in the 2-th layer. Here, we present the dimensions ranked in the top 14 and bottom 4. In the actual process, we iterate for 5 epochs on all samples on the training set to initialize AFFS. In Table 7, we present the changes in feature dimensions of different backbone features before and after performing AFFS
@@
Table 7: Feature dimension parameters of AFFS.
Method Training set CDF DFDCP Avg.
LAA-Net FF++(C23) +SBI 0.840 0.741 0.791
LAA-Net FF++(C23) +FIA-USA 0.875 0.782 0.829
Table 8: We report the Video-level AUC. The combination of our data augmentation method with other Sota detection algorithms.
@@ Appendix G Testing on more forgery techniques
In addition to several widely used forgery detection datasets mentioned in the paper, we also conducted experiments on other sota forgery techniques based on the DF40 dataset [87], where real samples were obtained from FF++ and fake samples were constructed using corresponding forgery techniques based on FF++ settings. As shown in the Table 9(a), our proposed forgery detection framework has achieved excellent detection results on the vast majority of forgery methods. Meanwhile, Table 9(b) shows the comparison between our method and other state-of-the-art methods on various forgery techniques in DIFF
@@
(a) We tested all forgery techniques provided on the DF40 dataset and reported Frame-Level AUC.
Method Train Set Test Subset
Real Fake Cycle CoDiff Imagic DiFace DCFace Dream SDXL_R FD_I LoRA Midj FD_T SDXL HPS
@@
[…]
Table 9: Test results on more forgery techniques.
@@ Appendix H Combining with SoTA detection method
In this section, we plan to combine our proposed FIA-USA data augmentation with other sota image-level detection methods to validate the effectiveness of our proposed data augmentation method. However, due to the limited availability of open-source image level detection methods, we can only provide the performance of our data augmentation method on LAA-Net [55] (CVPR 24) in Table 8. In the original paper, the model was trained on the FF++ dataset in the raw version, while here we trained on the C23 version.
Appendix I More visual examples

### v2

@@ 4.3 Ablation Study
The impact of method components. To systematically evaluate the contribution of each component in our framework, we conducted comprehensive ablation studies across seven benchmark datasets. As detailed in Table 6(a), our analysis focuses on three core innovations: (1) the FIA-USA augmentation mechanism, (2) Automatic Forgery-aware Feature Selection (AFFS), and (3) Region-aware Contrastive Regularization (RCR). We adopt video-level AUC for traditional deepfakes and frame-level AUC for generative deepfakes. The empirical results demonstrate progressive performance degradation when removing individual components, with the complete framework achieving optimal detection accuracy . The combination of FIA-USA, AFFS, and RCR yields optimal performance. As intended in our design, AFFS and RCR are crucial for fully leveraging the potential of FIA-USA. Removing any component degrades performance, with the absence of FIA-USA having the most significant impact (a 4.15% decrease). It is important to clarify that "w/o FIA-USA" specifically refers to using only the data augmentation proposed by SBI. Furthermore, RCR exerts a stronger influence on FIA-USA’s effectiveness compared to AFFS.
The ablation specifically targeting FIA-USA (components within it) reveals: Various data augmentation techniques contribute differently. SR and AE are shown to be effective for detecting generative forgeries. Conversely, MTMS appears to suppress performance on this specific forgery type. This aligns with our theoretical analysis: generative forgeries constitute global manipulations that produce fewer Face Inconsistency Artifacts (FIA) but primarily exhibit Up-Sampling Artifacts (USA). As seen from MTMS’s performance on traditional forgeries, increasing the diversity of FIA proves beneficial for detecting face swapping forgery.
Regarding the RCR ablation: The results align well with our chosen hyperparameters. They indicate that the PCR contributes substantially more to the model’s performance than the ICR. This observation is consistent with the hyperparameter sensitivity analysis presented in Table 6c: increasing the loss weight for PCR while decreasing that for ICR leads to significant performance gains. Conversely, increasing the ICR weight while reducing the PCR weight results in only marginal improvements.
The impact of model framework. Our architectural analysis in Table 6(b) systematically benchmarks detection performance across backbone networks, quantitatively comparing the effects of EfficientNet tan2019efficientnet , ResNet he2016deep , and different depth configurations on detection performance, demonstrates the compatibility of our method with various network frameworks, while also illustrating the critical impact of model parameter size and architectural design on detector performance. This highlights the ongoing importance of developing more robust and specialized deep learning architectures for counterfeit detection tasks.
Impact of hyperparameters. We also examined the effect of hyperparameters of the loss function on model performance, as detailed in the Table 6(c). Our analysis of the weights revealed that increasing the weight of $\lambda_{2}$ relative to the weights of cross entropy enhances model performance, whereas decreasing the weight of $\lambda_{3}$ relative to cross entropy also improves performance, which is consistent with the ablation experiment results for RCR components in Table 6(a). However, conversely, model performance will not be significantly improved.
@@ 5 Visualizations
Our GradCAM zhou2016learning analysis in Figure 4 demonstrates precise localization of facial forgery artifacts across both conventional and emerging generative architectures. Notably, the visualization framework not only pinpoints manipulation traces in traditional deepfakes but also effectively in state-of-the-art synthetic media, validating our method’s generalization capability.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported in part by Key Science & Technology Project of Anhui Province 202423l10050033, Zhejiang Provincial Natural Science Foundation of China under Grant (No. LQN25F020023).
@@ Appendix A Appendix
  • Sec. I: We discussed the limitations of this paper.
  • Sec. J: We discussed the impacts of this paper.
  • Sec. K: We provided more visual examples.
@@ Appendix D Details of FPN
As shown in the Figure 11, our FPN follows the framework proposed by [45], using the output features of the 2nd, 3rd, 4th, and 5th layers of the backbone to construct an FPN, In our experimental setup, the number of feature map channels is 128 and 196, corresponding to EfficientNet [66], and Resnet [29], respectively.
@@ E.2 AFFS
1) First, let’s consider multiple positive and negative sample pairs generated by FIA-USA, along with their corresponding masks $\{R_{n},F_{n},M_{n}\}_{n=1}^{N}$. The usual method is to directly construct an FPN using a set of features $f_{i}$ from a neural network $\phi$ as input, where $i$ represents the $i$-th layer of the $\phi$.
@@
$\mathcal{F}$ represents normalization and scaling operations, $f_{i}^{k}\in\mathbb{R}^{W_{i}*H_{i}},k\in[0,C_{i}]$, represents the feature map of the $k$-th dimension of the $f_{i}$ . During the training process, we only select the channels reserved for $f_{i}$ during the AFFS initialization process. After AFFS processing, the original features $f_{i}\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ will become $f_{i}^{\prime}\in\mathbb{R}^{W_{i}*H_{i}*m_{i}}$.
@@ E.3 Visual Explanation of AFFS
Here we provide a set of visual examples of AFFS in Figure 13. The experiment was conducted on the 2-th layer of EfficientNetB4. Unlike the actual situation, in order to enhance readability, we only calculated $L_{AFFS}$ for a single epoch on a single pair of samples. And based on this, sort the 32 dimensions of the features in the 2-th layer. Here, we present the dimensions ranked in the top 14 and bottom 4. In the actual process, we iterate for 5 epochs on all samples on the training set to initialize AFFS. In Table 7, we present the changes in feature dimensions of different backbone features before and after performing AFFS.
@@
Table 8: We report the Video-level AUC. The combination of our data augmentation method with other SOTA detection algorithms. $\dagger$} represents the method used in the original paper.
Method Training set CDFv2 DFDCP Avg.
LAA-Net SBI + EPPN $\dagger$} 0.840 0.741 0.791
LAA-Net FIA-USA + EFPN 0.875 0.782 0.829
Ours FIA-USA + FPN (AFFS) 0.901 0.861 0.881
Ours FIA-USA + FPN (AFFS) +RCR 0.941 0.866 0.904
@@ Appendix G Testing on more forgery techniques
In addition to several widely used forgery detection datasets mentioned in the paper, we also conducted experiments on other forgery techniques based on the DF40 dataset [80], where real samples were obtained from FF++ and fake samples were constructed using corresponding forgery techniques based on FF++ settings. As shown in the Table 9(a), our proposed forgery detection framework has achieved excellent detection results on the vast majority of forgery methods. Meanwhile, Table 9(b) shows the comparison between our method and other state-of-the-art methods on various forgery techniques in DIFF.
Table 9: Test results on more forgery techniques.
(a) We tested all forgery techniques provided on the DF40 dataset and reported Frame-Level AUC. All results with AUC greater than 80 are highlighted in bold, while those with AUC less than 80 but greater than 70 are underlined.
@@
(b) Comparison with universal deepfake detection methods using the Frame-Level AUC on the DiFF dataset. The notation $\dagger$}} indicates models that are designed for deepfake detection, while $\ddagger$}} signifies models intended for general generated detection. All experiments were trained on the c23 version of FF++. The best result is bolded.
Method Test Subset
Cycle CoDiff Imagic DiFace DCFace Dream SDXL_R FD_I LoRA Midj FD_T SDXL HPS
@@ Appendix H Combining with SoTA detection method
In this section, we aim to validate the effectiveness of our FIA-USA data augmentation by integrating it with state-of-the-art image-level detection methods. While our comprehensive evaluation currently focuses on LAA-Net [51] (CVPR 24) due to limited availability of open-source image-level detection implementations, Table 8 reveals significant performance improvements. Notably, the original LAA-Net implementation was trained on raw version FF++ data, whereas our experiments utilize the c23 version. Compared with the baseline implementation, our data augmentation strategy has significantly improved the performance of the original detector. To address requests for detailed comparisons with LAA-Net, our table shows two key findings: First, our method achieves a 7.5% performance advantage over LAA-Net when using identical data augmentation conditions. Second, even without employing RCR (Robust Context Refinement), our approach substantially outperforms LAA-Net’s EFPN component, demonstrating the inherent strength of our core methodology.
Appendix I Limitations
The proposed detection framework primarily focuses on image-level data augmentation and matching detection architectures. While we categorize image-level artifacts into FIA and USA, and have achieved satisfactory video-level detection performance, we acknowledge that video-level artifacts present more complex challenges - including temporal inconsistencies and audio-visual asynchrony - which fall outside the current research scope but represent our intended future research direction.
Second, although recent approaches increasingly incorporate large models’ zero-shot capabilities for deepfake detection, existing work has not yet explored the integration of data augmentation methods with such models due to practical constraints including variations in training methodologies and model scales. This does not mean that our method cannot be applied to large models, but currently we are still unfamiliar with large models and have not fully explored this part.
Regarding experimental validation, we mainly compared and analyzed our proposed method with previous data augmentation methods, which is consistent with the main contribution of our paper. Of course, we also compared state-of-the-art detection methods that are not data augmentation based. Following established practices in data augmentation research, we conducted comprehensive evaluations across four widely adopted network architectures to ensure fair comparison. However, we note that detection performance may vary across different network frameworks depending on their baseline performance and inherent compatibility with deepfake detection tasks. We cannot guarantee that all model frameworks will achieve excellent detection results after using our method.
Finally, while our method demonstrates effectiveness across a substantial majority of forgery methods (validated on over 58 different techniques), we acknowledge relatively weaker performance in specific edge cases. Nevertheless, the framework maintains robust detection capability for most common forgery approaches, achieving satisfactory overall performance that meets our research objectives.
Appendix J Broader impacts
The proposed framework offers positive societal impacts by enhancing detection of sophisticated face deepfakes, thereby mitigating disinformation campaigns, fake news, and fraudulent content that erode public trust in digital media. Additionally, it safeguards individual privacy by identifying forged facial manipulations that enable identity theft, unauthorized face swapping, and other privacy violations. The method could further support content authenticity initiatives through integration into social media platforms or news verification systems to prioritize legitimate visual content. However, potential negative impacts include the risk of adversaries reverse-engineering the detection framework to refine deepfake generation methods, potentially escalating the adversarial "arms race" between forgery creators and detectors. Furthermore, false positives in detection could lead to unintended censorship, where legitimate content is erroneously flagged as fake.
Appendix K More visual examples